Fill in the missing values in the following table. Assume that the value of the  MPC

LOADING...

does not change as real GDP changes and that there are zero taxes.   

​(Enter

all values as

integers.​)

Real GDP

​(Y)

Consumption

​(C)

Planned

Investment​ (I)

Government

Purchases​ (G)

Net

Exports​ (NX)

​$8 comma 0008,000

​$4 comma 0004,000

​$800

​$1 comma 2001,200

minus−​$200200

​$9 comma 0009,000

​$4 comma 5004,500

​$800

1 comma 2001,200

minus−​$200200

​$10 comma 00010,000

​$nothing

​$800

1 comma 2001,200

minus−​$200200

​$11 comma 00011,000

​$nothing

​$800

1 comma 2001,200

minus−​$200200

​$12 comma 00012,000

​$nothing

​$800

1 comma 2001,200

minus−​$200200
Answer #1

Y=C+I+G+NX

C=Y-(I+G+NX)

For third row

C=10,00010,000-(800+1,2001,200-200200)

C=988208200

For fourth row

C=11,00011,000-(800+1,2001,200-200200)

C=1088209200

For fifth row

C=12,00012,000-11,801,800

C=1188210200